# Connections between Combinations Without Specified Separations and Strongly Restricted Permutations, Compositions, and Bit Strings

### v1

@@ Abstract
Let $S_{n}$ and $S_{n,k}$ be, respectively, the number of subsets and $k$-subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ such that no two subset elements differ by an element of the set $\mathcal{Q}$. We prove a bijection between such $k$-subsets when $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ and permutations $\pi$ of $\mathbb{N}_{n+jm}$ with $k$ excedances satisfying $\pi(i)-i\in\{-m,0,jm\}$ for all $i\in\mathbb{N}_{n+jm}$. We also identify a bijection between another class of restricted permutation and the cases $\mathcal{Q}=\{1,q\}$. This bijection allows us to prove a conjectured recursion relation for the number of such permutations which corresponds to the case $\mathcal{Q}=\{1,4\}$. We also obtain recursion relations for $S_{n}$ and $S_{n,k}$ in the case $\mathcal{Q}=\{1,5\}$ by first obtaining related recursion relations for the numbers of closed walks of a given length on a particular class of directed pseudograph. We give some classes of $\mathcal{Q}$ for which $S_{n}$ is also the number of compositions of $n+q$ into a given set of allowed parts, where $q$ is the largest element of $\mathcal{Q}$. A bijection between the $k$-subsets for any $\mathcal{Q}$ and bit strings is also noted. Aided by this, an efficient algorithm for finding $S_{n}$ and $S_{n,k}$ is given. We also prove a bijection between $k$-subsets for a class of $\mathcal{Q}$ and the set representations of size $k$ of equivalence classes for the occurrence of a given length-($q+1$) subword within bit strings. We then formulate a straightforward procedure for obtaining the generating function for the number of such equivalence classes.
@@ 1 Introduction
Combinations without specified separations (or restricted combinations) refer to the subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ with the property that no two elements of a subset differ by an element of a set $\mathcal{Q}$, the elements of which are independent of $n$. We denote the total number of such subsets and the number of $k$-subsets (i.e., subsets of size $k$) by $S^{\mathcal{Q}}_{n}$ and $S^{\mathcal{Q}}_{n,k}$, respectively. We drop the superscripts when it is clear which set of disallowed differences they apply to. For any $\mathcal{Q}$, one always takes $S^{\mathcal{Q}}_{n<0}=S^{\mathcal{Q}}_{n,k<0}=S^{\mathcal{Q}}_{n<k,k}=0$. When there are no restrictions, we just have the number of ordinary combinations and so $S^{\{\}}_{n}=2^{n}$ (A000079 in the OEIS [20]) and $S^{\{\}}_{n,k}=\tbinom{n}{k}$ (A007318). It appears to be well-known that $S^{\{1,\ldots,q\}}_{n}=S^{\{1,\ldots,q\}}_{n-1}+S^{\{1,\ldots,q\}}_{n-q-1}+\delta_{n+q,0}$, where $\delta_{i,j}$ is 1 if $i=j$ and 0 otherwise, and $S^{\{1,\ldots,q\}}_{n,k}=\tbinom{n+q(1-k)}{k}$. These reduce to the classic results $S^{\{1\}}_{n}=F_{n+2}$, where $F_{n}$ is the $n$-th Fibonacci number (A000045, $F_{n}=F_{n-1}+F_{n-2}+\delta_{n,1}$, $F_{n<1}=0$), and $S^{\{1\}}_{n,k}=\tbinom{n+1-k}{k}$ (A011973) [12]. Expressions for $S_{n}^{\mathcal{Q}}$ and $S^{\mathcal{Q}}_{n,k}$ have also been obtained when $\mathcal{Q}=\{q\}$ [14, 19, 16, 15, 3] and $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ [18, 2].
@@
Any tiling of an $n$-board can be expressed as a tiling using metatiles which are gapless groupings of tiles that cannot be split into smaller gapless groupings by removing a gapless grouping of tiles [9]. The simplest metatiles when tiling a long enough board with squares and combs are a single square (denoted by $S$) and a $(w_{1},g_{1},\ldots,w_{t})$-comb with all the gaps filled by squares. We call the latter metatile a filled comb and it is denoted by $CS^{g}$, where $g=\sum_{i}g_{i}$. If the comb has one tooth it is just a $w_{1}$-omino (and also a metatile by itself) and $S$ and $C$ are then the only metatiles. Otherwise, as a consequence of the restricted-overlap nature of the tiling, a $C$ can always be placed so that its leftmost cell occupies the next available empty cell of the yet-to-be-completed metatile. If more empty cells within the yet-to-be-completed metatile result (which, if a $C$ is placed starting in the final cell of the final gap of the first comb on the board, will be the case iff $2w_{t}<q$ – see Lemma 3.1/2 in [1]) this can continue indefinitely resulting in there being an infinite number of possible metatiles.
@@
A systematic way to generate all possible metatiles and hence, in the simpler cases, obtain recursion relations for $B_{n}$ and $B_{n,k}$, is via a directed pseudograph (henceforth simply referred to as a digraph) in which each node has a label giving a bit-string representation of the remaining gaps (a 0 representing an empty cell, a 1 a filled one) starting at the next gap in the yet-to-be-completed metatile [10, 1] (see Figs. 1 and 2 for examples). The 0 node represents both the empty board and the completed metatile. Each arc represents the addition of a tile and has a label showing the type of tile and a subscript giving the increase in length of the incomplete metatile resulting from placing that tile, if that increase is not 0. The length increase from adding a $C$ is given by $q+1-d$, where $d$ is the number of binary digits in the node the arc starts from but with $d=0$ for the 0 node. A walk starting and ending at the 0 node but not visiting it in between represents a metatile and is denoted by the string of $C$ and $S$ corresponding to the arcs in the order visited. For example, in Fig. 1, the loop attached to the 0 node corresponds to the $S$ metatile and the cycle starting at the 0 node and then visiting the $10^{3}$ 1 node (an abbreviation for $0001$), the $10^{2}$ 1 node, and then the 01 node before returning to the 0 node corresponds to the $CS^{2}$ metatile. A cycle that does not include the 0 node is called an inner cycle (e.g., $SSC_{[4]}$ in Fig. 1 where including the length increment subscript quickly identifies which comb arc we are referring to). There are a finite number of possible metatiles iff the digraph lacks an inner cycle since a walk can traverse an inner cycle an arbitrary number of times before returning to the 0 node. If all the inner cycles have at least one node in common, that node is called a common node (e.g., the $01^{q-p}$ node is the common node in the Fig. 5 digraph; either the $10^{2}$ 1 or the 01 node can be chosen as the common node in the digraph in Fig. 2; Fig. 1 and Fig. 7 show digraphs that have inner cycles but no common node). For digraphs with a common node, as we review in §2, there are simple expressions for the recursion relations for numbers of tilings $B_{n}$ and $B_{n,k}$ in terms of lengths of various cycles in the digraph [10, 1]. If the digraph possesses inner cycles but no common node, as far as we are aware, there is only one other class of digraph (which has 3 inner cycles) where an analogous simple procedure has been found for obtaining the recursion relations [2].
@@
A systematic search of the OEIS for sequences $(S^{\mathcal{Q}}_{n})_{n\geq 0}$ for all $\mathcal{Q}$ with $q\leq 4$ showed that there are various connections between some types of restricted combinations and some types of restricted permutations. Relations to compositions and bit strings were also found. Here, in §3, we give bijections between between two classes of strongly restricted permutations and combinations with $\mathcal{Q}=\{1,q\}$ and $\mathcal{Q}=\{m,2m,\ldots,jm\}$. As a consequence of the former bijection, we can prove a conjectured recursion relation for the permutations corresponding to the $\mathcal{Q}=\{1,4\}$ case. In §2, we derive recursion relations for $B_{n}$ and $B_{n,k}$ for tilings with metatiles generated by a particular class of digraph with 4 inner cycles but no common node. This enables one to also obtain recursion relations for $S^{\{1,5\}}_{n}$ and $S^{\{1,5\}}_{n,k}$. In §4, although it is not all that surprising given the tiling interpretation of restricted combinations, for completeness, we also show that $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into certain sets of parts for some classes of $\mathcal{Q}$. An elementary connection between all instances of $\mathcal{Q}$ and bit strings is pointed out in §5. This forms the basis for an algorithm for efficiently calculating $S_{n}$ and $S_{n,k}$ for any given $\mathcal{Q}$ as detailed via a heavily annotated C program listed in the Appendix. In §5 we also give a bijection between $k$-subsets for certain $\mathcal{Q}$ and equivalence classes for the occurrence of certain subwords within bit strings. The bijection enables us to show that a generating function for the number of equivalence classes is always straightforward to obtain.
2 Recursion relations for numbers of walks on 4-inner-cycle digraphs with two errant loops
@@
We start by reviewing further terminology concerning digraphs. For a digraph possessing a common node, a walk from the 0 node to the common node followed by a walk from the common node back to the 0 node such that the common node is only visited once is called a common circuit (e.g., for either choice of common node, the digraph in Fig. 2 possesses a single common circuit, namely, $C_{[5]}S^{2}$). A cycle that includes the 0 node but not the common node is called an outer cycle (in Fig. 2, the $S$ loop attached to the 0 node is the only outer cycle).
@@
A form of the following theorem was originally stated in [10]. A more compact version, re-derived in [1], is reproduced here. We will use this theorem in §3 to obtain the $S^{\{1,4\}}_{n}$ generating function.
Theorem 2 (Theorem 4 in [1]).
@@
For a digraph possessing a common node, let $l_{\mathrm{o}i}$ be the length of the $i$-th outer cycle ($i=1,\ldots,N_{\rm o}$) and let $k_{\mathrm{o}i}$ be the number of combs it contains, let $L_{r}$ be the length of the $r$-th inner cycle ($r=1,\ldots,N$) and let $K_{r}$ be the number of combs it contains, and let $l_{\mathrm{c}i}$ be the length of the $i$-th common circuit ($i=1,\ldots,N_{\rm c}$) and let $k_{\mathrm{c}i}$ be the number of combs it contains. Then for all integers $n$ and $k$,
$B_{n}$ $=\delta_{n,0}+\sum_{r=1}^{N}(B_{n-L_{r}}-\delta_{n,L_{r}})+\sum_{i=1}^{N_{\rm o}}\biggl{(}B_{n-l_{\mathrm{o}i}}-\sum_{r=1}^{N}B_{n-l_{\mathrm{o}i}-L_{r}}\biggr{)}+\sum_{i=1}^{N_{\rm c}}B_{n-l_{\mathrm{c}i}},$
$B_{n,k}$ $=\delta_{n,0}\delta_{k,0}+\sum_{r=1}^{N}(B_{n-L_{r},k-K_{r}}-\delta_{n,L_{r}}\delta_{k,K_{r}})+\sum_{i=1}^{N_{\rm o}}\biggl{(}B_{n-l_{\mathrm{o}i},k-k_{\mathrm{o}i}}-\sum_{r=1}^{N}B_{n-l_{\mathrm{o}i}-L_{r},k-k_{\mathrm{o}i}-K_{r}}\biggr{)}$
$\qquad\mbox{}+\sum_{i=1}^{N_{\rm c}}B_{n-l_{\mathrm{c}i},k-k_{\mathrm{c}i}},$
where $B_{n<0}=B_{n,k<0}=B_{n<k,k}=0$.
@@
The following definitions are slightly modified versions of those given in earlier work on a class of digraph with inner cycles but no common node in order to incorporate the increased complexity of the type of digraphs we are about to consider [2]. For a digraph lacking a common node, we refer to any inner cycle that can be represented as a single arc linking a node to itself as an errant loop if the digraph would only have a common node $\mathcal{P}$ if all the errant loop arcs were removed. The node $\mathcal{P}$ is then referred to as a pseudo-common node. A node with an errant loop is called an errant loop node. Evidently, a particular errant loop node $\mathcal{E}$ cannot also be a pseudo-common node; if that were the case, removing all the other errant loop arcs except the errant loop at $\mathcal{E}$ would result in a digraph with a common node implying that the loop at $\mathcal{E}$ was never an errant loop in the first place. For a digraph with at least one errant loop, a common circuit is now defined as two concatenated simple paths from the 0 node to $\mathcal{P}$ and from $\mathcal{P}$ to the 0 node. We also need to modify the definition of an outer cycle: it is now a cycle that includes the 0 node but not $\mathcal{P}$.
The proof of Theorem 6 requires the following three results, the first of which is the $N=2$ case of Lemma 35 in [2].
Lemma 3.
For positive integers $j_{1},j_{2},j_{4}$ we have
$\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{2}+j_{4}-1}{j_{4}}\!=\!\binom{j_{1}+j_{2}-1}{j_{1}-1}\Biggl{(}\!\binom{j_{2}+j_{4}-1}{j_{4}}-\binom{j_{2}+j_{4}-2}{j_{4}-1}\!\Biggr{)}\\ +\binom{j_{1}+j_{2}-1}{j_{1}}\binom{j_{2}+j_{4}-2}{j_{4}}+\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{2}+j_{4}-2}{j_{4}-1}.$ (1)

### v2

@@ Abstract
Let $S_{n}$ and $S_{n,k}$ be, respectively, the number of subsets and $k$-subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ such that no two subset elements differ by an element of the set $\mathcal{Q}$, the largest element of which is $q$. We prove a bijection between such $k$-subsets when $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ and permutations $\pi$ of $\mathbb{N}_{n+jm}$ with $k$ excedances satisfying $\pi(i)-i\in\{-m,0,jm\}$ for all $i\in\mathbb{N}_{n+jm}$. We also identify a bijection between another class of restricted permutation and the cases $\mathcal{Q}=\{1,q\}$. This bijection allows us to prove a conjectured recursion relation for the number of such permutations which corresponds to the case $\mathcal{Q}=\{1,4\}$. We also obtain the generating function for $S_{n}$ in the case $\mathcal{Q}=\{1,5\}$ by first obtaining generating functions for the numbers of closed walks of a given length on a particular class of directed pseudograph. We give some classes of $\mathcal{Q}$ for which $S_{n}$ is also the number of compositions of $n+q$ into a given set of allowed parts. A bijection between the $k$-subsets for any $\mathcal{Q}$ and bit strings is also noted. Aided by this, an efficient algorithm for finding $S_{n}$ and $S_{n,k}$ is given. We also prove a bijection between $k$-subsets for a class of $\mathcal{Q}$ and the set representations of size $k$ of equivalence classes for the occurrence of a given length-($q+1$) subword within bit strings. We then formulate a straightforward procedure for obtaining the generating function for the number of such equivalence classes.
@@ 1 Introduction
Combinations without specified separations (or restricted combinations) refer to the subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ with the property that no two elements of a subset differ by an element of a set $\mathcal{Q}$, the elements of which are independent of $n$. We denote the total number of such subsets and the number of $k$-subsets (i.e., subsets of size $k$) by $S^{\mathcal{Q}}_{n}$ and $S^{\mathcal{Q}}_{n,k}$, respectively. We drop the superscript when it is clear which set $\mathcal{Q}$ of disallowed differences we are dealing with. For any $\mathcal{Q}$, one always takes $S^{\mathcal{Q}}_{n<0}=S^{\mathcal{Q}}_{n,k<0}=S^{\mathcal{Q}}_{n<k,k}=0$.
When there are no restrictions, we just have the number of ordinary combinations and so $S^{\{\}}_{n}=2^{n}$ (A000079 in the OEIS [20]) and $S^{\{\}}_{n,k}=\tbinom{n}{k}$ (A007318). It appears to be well-known that $S^{\{1,\ldots,q\}}_{n}=S^{\{1,\ldots,q\}}_{n-1}+S^{\{1,\ldots,q\}}_{n-q-1}+\delta_{n+q,0}$, where $\delta_{i,j}$ is 1 if $i=j$ and 0 otherwise, and $S^{\{1,\ldots,q\}}_{n,k}=\tbinom{n+q(1-k)}{k}$. These reduce to the classic results $S^{\{1\}}_{n}=F_{n+2}$, where $F_{n}$ is the $n$-th Fibonacci number (A000045, $F_{n}=F_{n-1}+F_{n-2}+\delta_{n,1}$, $F_{n<1}=0$), and $S^{\{1\}}_{n,k}=\tbinom{n+1-k}{k}$ (A011973) [11]. Expressions for $S_{n}^{\mathcal{Q}}$ and $S^{\mathcal{Q}}_{n,k}$ have also been obtained when $\mathcal{Q}=\{q\}$ [13, 19, 15, 14, 3] and $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ [17, 1].
@@
Any tiling of an $n$-board can be expressed as a tiling using metatiles which are gapless groupings of tiles that cannot be split into smaller gapless groupings by removing a gapless grouping of tiles [8]. The simplest metatiles when tiling a long enough board with squares and combs are a single square (denoted by $S$) and a $(w_{1},g_{1},\ldots,w_{t})$-comb with all the gaps filled by squares. We call the latter metatile a filled comb and it is denoted by $CS^{g}$, where $g=\sum_{i}g_{i}$. If the comb has one tooth it is just a $w_{1}$-omino (and also a metatile by itself) and $S$ and $C$ are then the only metatiles. Otherwise, as a consequence of the restricted-overlap nature of the tiling, a $C$ can always be placed so that its leftmost cell occupies the next available empty cell of the yet-to-be-completed metatile. If more empty cells within the yet-to-be-completed metatile result (which, if a $C$ is placed starting in the final cell of the final gap of the first comb on the board, will be the case iff $2w_{t}<q$ – see Lemma 2 in [2]), this can continue indefinitely and so there are an infinite number of possible metatiles.
@@
A systematic way to generate all metatiles, and thus obtain generating functions for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$, is to first construct a directed pseudograph (henceforth simply referred to as a digraph) in which each node has a label giving a bit-string representation of the remaining gaps and filled cells (a 0 representing an empty cell, a 1 a filled one) starting at the next gap in the yet-to-be-completed metatile and ending at the final filled cell [9, 2] (see Figs. 3 and 4 for examples). The 0 node represents both the empty board and the completed metatile. Each arc represents the addition of a tile and has a label showing the type of tile and a subscript giving the increase in length of the incomplete metatile resulting from placing that tile, if that increase is not 0. The length increase from adding a $C$ is given by $q+1-d$, where $d$ is the number of binary digits in the node the arc starts from but with $d=0$ for the 0 node. A walk starting and ending at the 0 node but not visiting it in between represents a metatile and is denoted by the string of $C$ and $S$ corresponding to the arcs in the order visited. For example, in Fig. 4, the loop attached to the 0 node corresponds to the $S$ metatile and the cycle starting at the 0 node and then visiting the $10^{3}$ 1 node (an abbreviation for $0001$), the $10^{2}$ 1 node, and then the 01 node before returning to the 0 node corresponds to the $CS^{2}$ (filled comb) metatile.
A cycle that does not include the 0 node is called an inner cycle (e.g., $SSC_{[4]}$ in Fig. 4, where including the length increment subscript quickly identifies which comb arc we are referring to). There are a finite number of possible metatiles iff the digraph lacks an inner cycle since a walk can traverse an inner cycle an arbitrary number of times before returning to the 0 node. If all the inner cycles have at least one node in common, that node is called a common node (e.g., the $01^{q-p}$ node is the common node in the Fig. 7 digraph; either the $10^{2}$ 1 or the 01 node can be chosen as the common node in the digraph in Fig. 3; Fig. 4 and Fig. 10 show digraphs that have inner cycles but no common node). For digraphs with a common node, there are simple expressions for the recursion relations for the numbers of tilings $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ in terms of lengths of various cycles in the digraph [9, 2] (we rederive these in the form of generating functions in §2). If the digraph possesses inner cycles but no common node, as far as we are aware, there is only one other class of digraph (which has 3 inner cycles) where an analogous simple procedure has been found for obtaining the recursion relations [1].
@@
A systematic search of the OEIS for sequences $(S^{\mathcal{Q}}_{n})_{n\geq 0}$ for all $\mathcal{Q}$ with $q\leq 4$ showed that there are various connections between some types of restricted combinations and some types of restricted permutations. Relations to compositions and bit strings were also found. Here, in §3, we give bijections between between two classes of strongly restricted permutations and combinations with $\mathcal{Q}=\{1,q\}$ and $\mathcal{Q}=\{m,2m,\ldots,jm\}$. As a consequence of the former bijection, we can prove a conjectured recursion relation for the permutations corresponding to the $\mathcal{Q}=\{1,4\}$ case. In §2, we derive generating functions for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ for tilings with metatiles generated by a particular class of digraph (which includes the 3 inner cycle class with no common node dealt with previously [1]) without a common node. This enables us to also obtain a generating function for $S^{\{1,5\}}_{n}$. In §4, although it is not all that surprising given the tiling interpretation of restricted combinations, for completeness, we also show that $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into certain sets of parts for some classes of $\mathcal{Q}$. An elementary connection between all instances of $\mathcal{Q}$ and bit strings is pointed out in §5. This forms the basis for an algorithm for efficiently calculating $S_{n}$ and $S_{n,k}$ for any given $\mathcal{Q}$ as detailed via a heavily annotated C program listed in the Appendix. In §5 we also give a bijection between $k$-subsets for certain $\mathcal{Q}$ and equivalence classes for the occurrence of certain subwords within bit strings. The bijection enables us to show that a generating function for the number of equivalence classes is always straightforward to obtain.
2 Generating functions for numbers of walks on digraphs with a pseudo-common node
@@
We start by reviewing further terminology concerning digraphs. For a digraph possessing a common node, a walk from the 0 node to the common node followed by a walk from the common node back to the 0 node (and so the common node is only visited once) is called a common circuit (e.g., for either choice of common node, the digraph in Fig. 3 possesses a single common circuit, namely, $C_{[5]}S^{2}$). A cycle that includes the 0 node but not the common node is called an outer cycle (e.g., in Fig. 3, the $S$ loop attached to the 0 node is the only outer cycle).
The proofs in this section require the following three lemmas [18]. The bivariate generating functions $G_{i}(x,y)$ we employ here are such that the coefficient of $x^{l}y^{k}$ of the generating function of a walk or set of walks between two nodes is the number of configurations of tiles that contribute $l$ to the overall length of the tiling and contain $k$ combs.
Figure 1: Schematic diagrams of connections between nodes corresponding to the construction of generating functions via (a) the parallel rule (Lemma 2), (b) the series rule (Lemma 3), and (c) the loop rule (Lemma 4).
Lemma 2 (The Parallel Rule).
If two nodes in a digraph are connected via $m$ alternative walks (or sets of walks), as shown in Fig. 1a, that have associated generating functions $G_{i}(x,y)$, where $i=1,\ldots,m$, then the combined generating function for all of those walks connecting the two nodes is $\sum_{i=1}^{m}G_{i}(x,y)$.
Lemma 3 (The Series Rule).
If two nodes in a digraph are connected via $m-1$ intermediate nodes which are connected by walks (or sets of walks), as shown in Fig. 1b, that have associated generating functions $G_{i}(x,y)$, where $i=1,\ldots,m$, then the combined generating function for all of those walks connecting the two nodes is $\prod_{i=1}^{m}G_{i}(x,y)$.
Lemma 4 (The Loop Rule).
If a node is connected to itself by a walk (or set of walks), as shown in Fig. 1c, with a corresponding generating function $G_{1}(x,y)$, then the overall generating function corresponding to the traversal of the loop walk (or walks) zero or more times is
$1+G_{1}(x,y)+G_{1}^{2}(x,y)+\cdots=\frac{1}{1-G_{1}(x,y)}.$
@@
A form of the following theorem was originally stated in [9], and a more compact version re-derived in [2]. Here, using the above three lemmas, we obtain the result in terms of a bivariate generating function. We will use this theorem in §3 to obtain the $S^{\{1,4\}}_{n}$ generating function.
Theorem 5.
@@
For a digraph possessing a common node, let ${l_{\mathrm{o}i}}$ be the length of the $i$-th outer cycle ($i=1,\ldots,N_{\rm o}$) and let ${k_{\mathrm{o}i}}$ be the number of combs it contains, let $L_{r}$ be the length of the $r$-th inner cycle ($r=1,\ldots,N$) and let $K_{r}$ be the number of combs it contains, and let ${l_{\mathrm{c}i}}$ be the length of the $i$-th common circuit ($i=1,\ldots,N_{\rm c}$) and let ${k_{\mathrm{c}i}}$ be the number of combs it contains. Then the generating function such that the coefficient of $x^{n}y^{k}$ thereof is the number of tilings an $n$-board that contain $k$ combs is given by
$G(x,y)=\frac{\displaystyle{1-\sum_{r=1}^{N}x^{L_{r}}y^{K_{r}}}}{\displaystyle{1-\sum_{r=1}^{N}x^{L_{r}}y^{K_{r}}-\sum_{i=1}^{N_{\rm o}}\biggl{(}x^{l_{\mathrm{o}i}}y^{k_{\mathrm{o}i}}-\sum_{r=1}^{N}x^{{l_{\mathrm{o}i}}+L_{r}}y^{{k_{\mathrm{o}i}}+K_{r}}\biggr{)}-\sum_{i=1}^{N_{\rm c}}x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}}}.$ (1)
Figure 2: Schematic diagrams of digraphs with (a) a common node $\mathcal{C}$ and (b) a pseudo-common node $\mathcal{P}$ to which Theorem 5 and Theorem 6 apply, respectively. In these diagrams, each arc, with the exception of each errant loop (each arc connecting an $\mathcal{E}_{i}$ node to itself), potentially represents a number of alternative arcs in the specific instance of a digraph. Some nodes and arcs can be absent.
Proof.
The generating functions corresponding to a single traversal of the $i$-th outer cycle (represented by the loop attached to the 0 node in Fig. 2a), the $i$-th common circuit (represented by the path from the 0 node to the $\mathcal{C}$ node and back or the path $0\to\mathcal{C}\to\mathcal{R}\to 0$), and the $r$-th inner cycle ($\mathcal{C}\to\mathcal{R}\to\mathcal{C}$) are, respectively, $x^{l_{\mathrm{o}i}}y^{k_{\mathrm{o}i}}$, $x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}$, and $x^{L_{r}}y^{K_{r}}$. The $N$ inner cycles are alternative paths from the common node back to itself and so, from the parallel and loop rules, the generating function corresponding to their traversal (starting from and returning to the common node) in any order zero or more times is $1/(1-\sum_{i=1}^{N}x^{L_{r}}y^{K_{r}})$. Using the series rule, we multiply this by $x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}$ to obtain the generating function for the path along the $i$-th common circuit that includes an arbitrary number of trips around the inner cycles. We now apply the parallel rule to all such cases and to all outer cycles, followed by the loop rule since these can be executed any number of times. This gives
$G(x,y)=\frac{\displaystyle{1}}{\displaystyle{1-\sum_{i=1}^{N_{\rm o}}x^{l_{\mathrm{o}i}}y^{k_{\mathrm{o}i}}-\frac{\displaystyle{\sum_{i=1}^{N_{\rm c}}x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}}}{\displaystyle{1-\sum_{r=1}^{N}x^{L_{r}}y^{K_{r}}}}}},$
which yields (1) on rearranging. ∎
@@
The following definitions are slightly modified versions of those given in earlier work on a class of digraph with inner cycles but no common node in order to incorporate the increased complexity of the type of digraphs we are about to consider [1]. For a digraph lacking a common node, we refer to any inner cycle that can be represented as a single arc linking a node to itself as an errant loop if the digraph would only have a common node $\mathcal{P}$ if all the errant loop arcs were removed. The node $\mathcal{P}$ is then referred to as a pseudo-common node. A node with an errant loop is called an errant loop node. Evidently, a particular errant loop node $\mathcal{E}$ cannot also be a pseudo-common node; if that were the case, removing all the other errant loop arcs except the errant loop at $\mathcal{E}$ would result in a digraph with a common node implying that the loop at $\mathcal{E}$ was never an errant loop in the first place. For a digraph with at least one errant loop, a common circuit is now defined as two concatenated simple paths from the 0 node to $\mathcal{P}$ and from $\mathcal{P}$ to the 0 node. We also need to modify the definition of an outer cycle: it is now a cycle that includes the 0 node but not $\mathcal{P}$. Inner cycles (other than errant loops), outer cycles, and common circuits are said to be plain if they do not contain any errant loop nodes and non-plain otherwise.
Theorem 6.
If a digraph has plain outer cycles of lengths ${l_{\mathrm{o}i}}$ that contain ${k_{\mathrm{o}i}}$ combs ($i=1\ldots,N_{\rm o}$), non-plain outer cycles of lengths ${\tilde{l}_{\mathrm{o}i}}$ that contain ${\tilde{k}_{\mathrm{o}i}}$ combs with associated errant loops having lengths ${l_{\mathrm{oe}i}}$ and ${k_{\mathrm{oe}i}}$ combs ($i=1\ldots,\tilde{N}_{\rm o}$), plain inner cycles (excluding the errant loops) of lengths $L_{r}$ that contain $K_{r}$ combs ($r=1,\ldots,N$), non-plain inner cycles of lengths $\tilde{L}_{r}$ that contain $\tilde{K}_{r}$ combs with associated errant loops having lengths $l_{\mathrm{e}r}$ and ${k_{\mathrm{e}r}}$ combs ($r=1,\ldots,\tilde{N}$), plain common circuits of lengths ${l_{\mathrm{c}i}}$ that contain ${k_{\mathrm{c}i}}$ combs ($i=1,\ldots,N_{\rm c}$), non-plain common circuits of lengths $\tilde{l}_{\mathrm{c}i}$ that contain ${\tilde{k}_{\mathrm{c}i}}$ combs with associated errant loops having lengths $l_{\mathrm{ce}i}$ and ${k_{\mathrm{ce}i}}$ combs ($i=1,\ldots,\tilde{N}_{\rm c}$), and all non-plain cycles and circuits contain exactly one errant node, then the generating function whose coefficient of $x^{n}y^{k}$ gives the number of tilings of an $n$-board that use $k$ combs is given by
$G(x,y)=\frac{\displaystyle{1}}{\displaystyle{1-\sum_{i=1}^{N_{\rm o}}x^{l_{\mathrm{o}i}}y^{k_{\mathrm{o}i}}-\sum_{i=1}^{\tilde{N}_{\rm o}}\frac{\displaystyle{x^{\tilde{l}_{\mathrm{o}i}}y^{\tilde{k}_{\mathrm{o}i}}}}{\displaystyle{1-x^{l_{\mathrm{oe}i}}y^{k_{\mathrm{oe}i}}}}-\frac{\displaystyle{\sum_{i=1}^{N_{\rm c}}x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}+\sum_{i=1}^{\tilde{N}_{\rm c}}\frac{\displaystyle{x^{\tilde{l}_{\mathrm{c}i}}y^{\tilde{k}_{\mathrm{c}i}}}}{\displaystyle{1-x^{l}_{\mathrm{ce}i}y^{k_{\mathrm{ce}i}}}}}}{\displaystyle{1-\sum_{r=1}^{N}x^{L_{r}}y^{K_{r}}-\sum_{r=1}^{\tilde{N}}\frac{\displaystyle{x^{\tilde{L}_{r}}y^{\tilde{K}_{r}}}}{\displaystyle{1-x^{l_{\mathrm{e}r}}y^{k_{\mathrm{e}r}}}}}}}}.$ (2)
Proof.
The proof is similar to that of Theorem 5; we simply have more possible types of path. The generating functions corresponding to a single traversal of the $i$-th plain outer cycle, the $i$-th plain common circuit, and the $r$-th plain inner cycle are, respectively, $x^{l_{\mathrm{o}i}}y^{k_{\mathrm{o}i}}$, $x^{l_{\mathrm{c}i}}y^{k_{\mathrm{c}i}}$, and $x^{L_{r}}y^{K_{r}}$. On application of the series and loop rules, one sees that the generating functions corresponding to a single traversal of the $i$-th non-plain outer cycle, the $i$-th non-plain common circuit, and the $r$-th non-plain inner cycle in addition to an arbitrary number of traversals of their errant loops are, respectively, $x^{\tilde{l}_{\mathrm{o}i}}y^{\tilde{k}_{\mathrm{o}i}}/(1-x^{l_{\mathrm{oe}i}}y^{k_{\mathrm{oe}i}})$, $x^{\tilde{l}_{\mathrm{c}i}}y^{\tilde{k}_{\mathrm{c}i}}/(1-x^{l_{\mathrm{ce}i}}y^{k_{\mathrm{ce}i}})$, and $x^{L_{r}}y^{K_{r}}/(1-x^{l_{\mathrm{e}r}}y^{k_{\mathrm{e}r}})$. In Fig. 2b these are represented, respectively, by the paths $0\to\mathcal{E}_{0}\to 0$, $0\to\mathcal{P}\to\mathcal{E}_{1}\to 0$ or $0\to\mathcal{P}\to\mathcal{E}_{2}\to\mathcal{R}\to 0$, and $\mathcal{P}\to\mathcal{E}_{2}\to\mathcal{R}\to\mathcal{P}$, where, on reaching $\mathcal{E}_{i}$, its loop can be executed an arbitrary number of times. It should now be apparent where the terms in the denominator of (2) arise from. The first two sums are from the plain and non-plain outer cycles. The numerator of the final term is the generating function for all the common circuits (including arbitrary number of traversals of the errant loops in the case of the non-plain ones). On reaching $\mathcal{P}$ via the first part of a common circuit, the inner cycles (and their associated errant loops in the case of the non-plain ones) can be executed in any order any number of times, which is what the denominator of the final term corresponds to (by application of all three rules). ∎